Determine a probability

assign a probability to the indicated event on the basis of information provided. indicate the technique you used: intuition, relative frequency, or the formula for equally likely outcomes.

a. a random sample of 500 students at Hudson college were surveyed and it was determined that 375 wear glasses or contact lenses. Estimate the probability that a Hudson college student selected at random wears corrective lenses.

b. the friends at the library host a fundraising barbeque. George is on the cleanup committeee. there are four members on this committee, and they draw lots to see who will clean the grills, assuming that each member is equally likely to be drawn. what is the probability that George will be assigned the grill cleanup job. P (George ) = # of favorable outcomes and total # of outcomes = 1/4 =0.25

c. Joanna photographs whales for sea life adventure fims. on her next expedetion, she is to film blue whales feeding. based on her knowledge of the habits of blue whales, she is almost certain she will be successful. what specific number do you suppose she estimates for the probability of success
Answer #1

a)

Probability = 375/500=0.75

b)

P(george) = 1/4 = 0.25

c)

0.90 < P(success) <1.00